A 20.0 g ball of clay traveling east at 4.00 m/s collides and sticks together with a 50.0 g ball of clay traveling north at 6.00 m/s. What is the speed of the resulting ball of clay?
Answer #1

By Conservation of momentum

Pi=Pf

m1V1 +m2V2 =(m1+m2)*V

20*(4i) +50*(6j) =(20+50)*V

V=1.143 i + 4.286 j

So speed of the resulting ball of clay is

|V| =sqrt[1.1432+4.2862]

|V|=4.44 m/s